OpenStudy (anonymous):

Nathaniel is 72-years-old and a heart attack survivor. His doctor tests his heart regularly with an electrocardiogram to ensure his heart has not worsened. Through a coinsurance agreement, his health insurance policy pays 65% of the cost for tests 1 – 4, 70% of the cost for tests 5 – 8, 75% of the cost for tests 9 – 12, and 80% of the cost for scans 13 – 16. Nathaniel must pay for any tests beyond the 16th. If each test costs $140 and he receives 22 each year, what is Nathaniel’s out-of-pocket expense? @ganeshie8

ganeshie8 (ganeshie8):

Total cost of tests = 22 * 140 = ?

OpenStudy (anonymous):

3080

ganeshie8 (ganeshie8):

next, calculate how much insurance pays him : 140( 4*0.65 + 4*0.7 + 4*0.75 + 4*0.8 ) = ?

OpenStudy (anonymous):

1624

ganeshie8 (ganeshie8):

yes, subtract both

OpenStudy (anonymous):

3080 - 1642 = 1456

ganeshie8 (ganeshie8):

out of pocket expense = "total cost" - "money paid by insurance"

ganeshie8 (ganeshie8):

1456 is right !

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

thank u !

ganeshie8 (ganeshie8):

u wlc :)